OpenStudy (anonymous):

Suppose, ABCD is a rectangle and that E is a point on CD. Let x be the area of AED, y be the area of BCE and z be the area of ABE and suppose y^2 = xz. What is the value of DE/EC ?

ganeshie8 (ganeshie8):

question looks incomplete ?

OpenStudy (anonymous):

yes diagram

OpenStudy (anonymous):

it takes so much time to load openstudy page.so coudnt able to post properly

OpenStudy (anonymous):

e is point on cd.what is the value of de/bc?

ganeshie8 (ganeshie8):

I have modified the question, see if it looks correct now ?

ganeshie8 (ganeshie8):

\[\large y = \dfrac{EC\times BC}{2}\] \[\large x = \dfrac{DE\times BC}{2}\] \[\large z = \dfrac{(DE+EC)\times BC}{2}\]

ganeshie8 (ganeshie8):

plug them in the given relation : \(\large y^2 = xz\)

ganeshie8 (ganeshie8):

\[\large \left(\dfrac{EC\times BC}{2}\right)^2 = \left(\dfrac{DE\times BC}{2}\right)\left(\dfrac{(DE+EC)\times BC}{2}\right)\]

ganeshie8 (ganeshie8):

canceling BC^2/4 both sides gives you : \[\large EC^2 = DE(DE+EC)\]

ganeshie8 (ganeshie8):

\[\large EC^2 = DE^2+DE\times EC\] \[\large 1 = \dfrac{DE^2}{EC^2}+\dfrac{DE}{ EC}\] \[\large 1 = \left(\dfrac{DE}{EC}\right)^2+\dfrac{DE}{ EC}\]

ganeshie8 (ganeshie8):

its a quadratic, you can solve it using quadratic formula

ganeshie8 (ganeshie8):

\[\large 1 = x^2 + x\] \[\large x^2 + x-1=0\] \[\large x = \dfrac{-1\pm \sqrt{1^2 - 4(1)(-1)}}{2(1)}\] \[\large = \dfrac{-1\pm \sqrt{5}}{2}\]

OpenStudy (anonymous):

hey how u modified my question ?

ganeshie8 (ganeshie8):

moderators can modify/delete questions :)

OpenStudy (anonymous):

hmmmmmmmmmmm.........great

ganeshie8 (ganeshie8):

since \(\large \dfrac{DE}{EC}\) is a ratio of distances, it cannot be negative, so the ratio has to be : \(\large \dfrac{-1+\sqrt{5}}{2}\)

ganeshie8 (ganeshie8):

see if that makes more or less sense..

OpenStudy (anonymous):

yes.how u are saying ec*bc/2

ganeshie8 (ganeshie8):

\(y\) is the area of right most triangle : base = \(EC\) height = \(BC\)

OpenStudy (anonymous):

u rocked!

OpenStudy (anonymous):

r u a maths professor.

ganeshie8 (ganeshie8):

lol no, that title was just a joke :)

OpenStudy (anonymous):

why u are giving medal to me.i am the questionnaire.:-)

ganeshie8 (ganeshie8):

cos i notice that you always participate actively in discussions about ur question and ask back good good questions :) thats the reason i gave u full solution

OpenStudy (anonymous):

no,no really i asked ru.or u should loved the maths most than any other.please tell me some tips to get threw.or suggest me some books

OpenStudy (anonymous):

ughh..my page is not loading.thats why it takes time for my reply

ganeshie8 (ganeshie8):

I think, taking time to understand fully below things help you in approaching any problem in geometry confidently : 1) how/why similar triangle ratios work (AA/SAS/SSS/AAS congruence and similarity applications) 2) making sense of all the area/volume formulas (triangles, six types of quadrilaterals : [parallelogram, rhombus, rectangle, square, trapezoid, kite], prisms, pyramids etc) 3) full understanding of four centers of triangle : centroid, incenter, circumcenter, orthocenter

ganeshie8 (ganeshie8):

50% of medium-to-hard problems in geometry involve dealing with similar/congruent triangles !

ganeshie8 (ganeshie8):

circles and inscribed polygons is also an interesting topic worth spending time

OpenStudy (anonymous):

ya came through these topic formulae,definition n all.but still

ganeshie8 (ganeshie8):

that `but still...` will disappear if we try to derive everything on our own and stop trusting all the formulas at their face value :P

ganeshie8 (ganeshie8):

when u have time, try to derive below well known formula for area of triangle :  \(\large \text{Area} = \dfrac{b*h}{2}\)